# mode = titleGraph Filtering for Multimodal Recommendation

### v2

@@ Abstract
Multimodal recommender systems improve the performance of canonical recommender systems with no item features by utilizing diverse content types such as text, images, and videos, while alleviating inherent sparsity of user–item interactions and accelerating user engagement. However, current neural network-based models often incur significant computational overhead due to the complex training process required to learn and integrate information from multiple modalities. To address this challenge, we propose MultiModal-Graph Filtering (MM-GF), a training-free method grounded in graph filtering (GF) for efficient and accurate multimodal recommendations. Specifically, MM-GF first constructs multiple similarity graphs for two distinct modalities as well as user–item interaction data. Then, MM-GF optimally fuses these multimodal signals using a polynomial graph filter that allows for precise control of the frequency response by adjusting frequency bounds. Furthermore, the filter coefficients are treated as hyperparameters, enabling flexible and data-driven adaptation. Extensive experiments on real-world benchmark datasets demonstrate that MM-GF not only improves recommendation accuracy by up to 22.25% compared to the best competitor but also dramatically reduces computational costs by achieving the runtime of less than 10 seconds.
@@ 1.1 Background and Motivation
Recently, multimodal recommender systems (MRSs) have garnered significant attention due to their ability to accommodate diverse item information from multiple modalities for enhanced recommendation performance. Compared to canonical recommender systems Ranjbar et al. (2015); Valcarce et al. (2019); He et al. (2020) with no item features (referred to as single-modal recommender systems), MRSs can capture and leverage precise item information (e.g., textual and/or visual features) for recommendations, thereby enhancing the overall capabilities of recommender systems He and McAuley (2016b); Wei et al. (2020). Notably, while single-modal recommender systems often struggle with sparsity of user–item interactions, MRSs can overcome this limitation by utilizing multimodal features of items. Due to these advantages, MRSs Zhang et al. (2021); Zhou et al. (2023b); Zhou and Shen (2023); Yu et al. (2023) are shown to substantially outperform single-modal recommendation methods based on collaborative filtering that rely solely on historical user–item interactions.
@@
Various MRSs have been developed to improve recommendation performance. In particular, thanks to the expressive capability via message passing in graph convolutional networks (GCNs) Kipf and Welling (2017), attention has been paid to GCN-based MRSs Wei et al. (2020); Zhang et al. (2021); Zhou et al. (2023b); Zhou and Shen (2023); Yu et al. (2023). For example, prior studies on MRSs learned GCNs separately to process different modalities, distinct from user–item interactions Wei et al. (2020); Zhang et al. (2021); Kim et al. (2024), or constructed an item–item similarity graph based on multimodal information (i.e., multimodal features of items) and then applied GCNs to process the similarity graph Zhang et al. (2021).
@@
Figure 1: Training time comparison of various MRSs under different degrees of modality information on the Baby dataset. Here, the end-to-end processing time is measured for our method that does not need any training process.
@@
On one hand, user preferences tend to shift quickly under the influence of trends, personal situations, and exposure to new content Ju et al. (2015); Pereira et al. (2018); hence, recommender systems need to be flexible to adapt to such dynamic preferences. Especially in environments where the training and inference speed is crucial, model runtime can become a significant bottleneck. In this context, incorporating additional information (i.e., multimodal features of items) into modeling significantly increases computational overhead for GCN-based MRSs. Consequently, MRSs learn multimodal information through GCNs, naturally leading to escalation of training time for GCNs with the increased number of modalities. Figure 1 illustrates how the training time (in seconds) of various MRSs behaves according to different degrees of modality information on the Baby dataset (one of well-known benchmark datasets for MRSs). As shown in the figure, the training time of GRCN Wei et al. (2020) and LATTICE Zhang et al. (2021) for processing all multimodal information is considerably high, which signifies the significant computational cost of GCN-based models when such data are integrated. Thus, although using GCN-based MRSs (e.g., Wei et al. (2020); Zhang et al. (2021)) is promising, we face practical challenges in computational complexity especially for the case when the speed at which recommender systems update their models is of paramount importance due to the dynamics of user preferences.
@@
On the other hand, recent studies on single-modal recommender systems Shen et al. (2021); Xia et al. (2022); Liu et al. (2023); Park et al. (2024) have adopted the training-free graph filtering (GF) (also known as graph signal processing) mechanism due to its simplicity and effectiveness. Such GF is a fundamental operation enabling the manipulation of signals defined over the nodes based on the underlying graph topology Defferrard et al. (2016); Isufi et al. (2024). A pioneering study on training-free GF is GF-CF Shen et al. (2021), which presented a novel approach to constructing an item–item similarity graph and applying GF that does not necessitate model training to enhance recommendation performance. Subsequent studies Liu et al. (2023); Xia et al. (2022) have been introduced to further enhance the effectiveness of GF-based recommender systems. Additionally, recent studies Park et al. (2024, 2025); Kim et al. (2025) proposed a powerful normalization scheme that enables adjustable influence between users and items, while introducing tunable hyperparameters to mitigate over-smoothing and under-smoothing and employing polynomial graph filters that avoid explicit eigendecomposition.
@@
Despite these advances, the GF-based recommendation methods in Park et al. (2024, 2025); Kim et al. (2025) present a key limitation from the perspective of frequency analysis. Normalization techniques in Shen et al. (2021); Xia et al. (2022); Liu et al. (2023) ensure that all eigenvalues of the underlying item–item similarity graph lie within the interval $[0,1]$, providing a stable theoretical foundation for designing polynomial filters in the frequency domain. In contrast, the normalization and additional adjustment process in Park et al. (2024, 2025); Kim et al. (2025) does not enforce this spectral constraint, allowing eigenvalues to extend beyond the unit interval, which distorts the intended frequency response. This effect is magnified on multimodal information due to heterogeneous characteristics, resulting in greater spectral instability than the case of user–item interactions. Additionally, such an effect can be amplified with higher-order polynomials, disproportionately emphasizing components tied to large-magnitude eigenvalues and further degrading frequency control. The aggregation of multiple out-of-band components generates spurious, highly distorted signals. These limitations underscore the need for a rigorous and sophisticated polynomial graph filtering strategy tailored for recommendation systems. Moreover, their polynomial filters in Park et al. (2024, 2025); Kim et al. (2025) were analytically derived from a fixed frequency response function, which lacks flexibility. This limitation underscores the need for a more adjustable and theoretically grounded filtering strategy for recommender systems.
@@ 1.2 Main Contributions
$\lambda_{min}$ Minimum eigenvalue of $\Lambda$
$\lambda_{max}$ Maximum eigenvalue of $\Lambda$
@@ 2 Related work
GF-based recommendation. In the context of GF, GCN Kipf and Welling (2017) is viewed as a parameterized convolutional filter for graphs. A notable GCN-based recommendation method is NGCF Wang et al. (2019), which was proposed to learn suitable low-pass filters (LPFs) while capturing high-order collaborative signals present in user–item interactions. LightGCN He et al. (2020) demonstrated strong performance by simplifying NGCF, removing both linear transformations and non-linear activations from its GCN layers. By bridging the gap between LightGCN and GF approaches, GF-CF Xu et al. (2019) was introduced, offering satisfactory recommendation accuracy with minimal computational costs due to its training-free design and a closed-form solution for the infinite-dimensional LightGCN. As a follow-up study, PGSP Liu et al. (2023) employed a mixed-frequency filter that integrates a linear LPF with an ideal LPF. Furthermore, in Turbo-CF Park et al. (2024), polynomial LPFs were designed to retain low-frequency signals without an ideal LPF that requires costly matrix decompositions.
Multimodal recommendation. Studies on MRSs have been actively conducted to boost the performance of recommender systems by leveraging information of multimodal features of items (i.e., textual and visual features). Initially, VBPR He and McAuley (2016b) extended Bayesian personalized ranking (BPR) loss Rendle et al. (2012) by utilizing visual features of items. Similarly as in other graph-based training models, GCNs have also gained attention due to their ability to seamlessly integrate multimodal information. For example, GRCN Wei et al. (2020) was presented alongside GCNs that construct a refined user–item bipartite graph so as to identify false-positive interactions by utilizing multimodal information. LATTICE Zhang et al. (2021) applied graph convolution operations to capture high-order item–item relationships and integrate them. BM3 Zhou et al. (2023b) presented a self-supervised learning framework, bootstrapping latent representations of both user–item interactions and multimodal features, thus providing a simple yet efficient approach to recommendation. FREEDOM Zhou and Shen (2023) extended LATTICE Zhang et al. (2021) by simplifying the construction of item–item similarity graphs and incorporating a degree-sensitive edge pruning method. In MGCN Yu et al. (2023), noisy features was purified, the purified modality features of items and behavior features were enriched in separate views, and a behavior-aware fuser was designed to predict user preferences. LGMRec Guo et al. (2024) exploited multimodal information through the construction of a hypergraph, which enables the joint modeling of various user preferences. PGL Yu et al. (2025) enhanced user preference modeling by extracting principal subgraphs from user–item interactions and performing message passing to jointly capture global and local information for principal subgraphs.
@@ 3.1 Notion of GF
We provide fundamental principles of GF (or equivalently, graph signal processing) Ortega et al. (2018); Shen et al. (2021); Park et al. (2024). First, we consider an undirected graph $\mathcal{G=(V,E)}$, which is represented by an adjacency matrix $A$ that indicates the presence or absence of edges between nodes. The Laplacian matrix $L$ of $\mathcal{G}$ is defined as L=DAL=D-A, where $D$ is the degree matrix of $A$ Grone and Merris (1994). Meanwhile, a graph signal on $\mathcal{G}$ is expressed as $\mathbf{x}\in\mathbb{R}^{|\mathcal{V}|}$, where each $x_{i}$ indicates the signal strength at the corresponding node $i$ in $\bf x$}. The smoothness of a graph signal $\mathbf{x}$ can be mathematically measured by
@@ 4 Methodology
In this section, we elaborate on the proposed MM-GF method as well as our research motivation.
@@ 4.1 Motivation and Challenges
Conventional training-free GF-based recommendation methods Shen et al. (2021); Liu et al. (2023) start by constructing a graph structure, where nodes represent items and edges correspond to item–item similarities. The graph construction process is formulated as follows:
@@
Recently, Park et al. (2024, 2025); Kim et al. (2025) utilized asymmetric normalization on $R$ to regularize the popularity of users or items before calculating $\tilde{P}$, which is formulated as follows:
@@
However, as critical issues, introducing two hyperparameters (i.e., $\alpha$ and $s$) may alter the spectral characteristics of the underlying graph. Specifically, the work in Shen et al. (2021) proved that all the eigenvalues of $\tilde{P}=\tilde{R}^{T}\tilde{R}$ where $\tilde{R}=D_{r}^{-0.5}RD_{c}^{-0.5}$ are in $[0,1]$. Nonetheless, as long as parameters $\alpha$ and $s$ are concerned, one cannot always assure that $\bar{P}$ in Eqs. (5) and (6) satisfies the aforementioned condition in Section 1.1 (i.e., the unit interval of eigenvalues) when $\alpha\neq 0.5$ and/or $s\neq 1$. In Figure 2, we present two cases: Case 1) $\alpha=0.5$ and $s=1$, corresponding to symmetric normalization, and Case 2) $\alpha=0.7$ and $s=0.6$, corresponding to asymmetric normalization with an additional adjustment process. The observations from this figure pose practical challenges below.
@@
As observed in Figure 2, in general, if we naïvely adopt the polynomial LPFs used in Park et al. (2024, 2025); Kim et al. (2025), then several issues arise. First, the filtered eigenvalues fall outside the range of $[0,1]$. Symmetric normalization in Shen et al. (2021); Liu et al. (2023); Xia et al. (2022) typically follow that the spectrum lies within this interval, which was proved by Shen et al. (2021). When eigenvalues are mapped out of this range, the polynomial filter may yield inaccurate frequency responses, deviating from its designed frequency characteristics. Moreover, as the polynomial order increases, the filter response for large-magnitude eigenvalues can grow dramatically; this results in a disproportionate emphasis on certain frequency components, thereby degrading the model performance. For example, consider a polynomial LPF $h(\lambda)=1-\lambda^{2}$. In a bounded frequency range of $[0,1]$, filtered values close to 0 correspond to high-frequency components, whereas those close to 1 correspond to low-frequency components. However, if the spectrum contains eigenvalues such as $\lambda_{1}\simeq-1$ and $\lambda_{2}\simeq 14$, as shown in Figure 2(b), then their corresponding filter responses become $h(\lambda_{1})\simeq 0$ and $h(\lambda_{2})\simeq-195$, respectively. This phenomenon results in signals being assigned to anomalous frequencies instead of the targeted ones. Furthermore, this problem becomes particularly critical for multimodal information, where signals are more susceptible to distortion compared to the case of user–-item interactions. For example, if the spectrum contains an eigenvalue $\lambda_{3}\simeq 70$, as illustrated in Figures 2(d) and 2(f), then its corresponding filter response becomes $h(\lambda_{3})\simeq-4899$, leading to far more severe distortion than the case of user-–item interactions. Ultimately, since MRSs require integrating signals across multiple modalities, aggregating such severely distorted signals can undermine the overall reliability of the recommendation process.
@@
Addressing a different issue, while recent GF models Park et al. (2024, 2025); Kim et al. (2025) demonstrated that a polynomial graph filter can be implemented as the form of $\sum_{k=1}^{K}a_{k}\bar{P}^{k}$, their empirical application was confined only to specific filters. This limited scope highlights a gap in the broader applicability of the GF methodology. Moreover, it is of paramount importance to effectively aggregate information across multiple modalities for accurate multimodal recommendations. In light of these challenges, a key question arises: "How can we design an efficient and effective GF method for multimodal recommendations by not only addressing the issue of filter design but also maximally exploiting heterogeneous characteristics across modalities?" To answer this question, we will outline the proposed MM-GF method tailored for multimodal recommendations in the following subsection.
Figure 3: The schematic overview of $M$-GF.

### v3

@@ Abstract
Multimodal recommender systems improve the performance of canonical recommender systems with no item features by utilizing diverse content types such as text, images, and videos, while alleviating inherent sparsity of user–item interactions and accelerating user engagement. However, current neural network-based models often incur significant computational overhead due to the complex training process required to learn and integrate information from multiple modalities. To address this challenge, we propose a training-free multimodal recommendation method grounded in graph filtering, designed for multimodal recommendation systems to achieve efficient and accurate recommendation. Specifically, the proposed method first constructs multiple similarity graphs for two distinct modalities as well as user–item interaction data. Then, it optimally fuses these multimodal signals using a polynomial graph filter that allows for precise control of the frequency response by adjusting frequency bounds. Furthermore, the filter coefficients are treated as hyperparameters, enabling flexible and data-driven adaptation. Extensive experiments on real-world benchmark datasets demonstrate that the proposed method not only improves recommendation accuracy by up to 22.25% compared to the best competitor but also dramatically reduces computational costs by achieving the runtime of less than 10 seconds.
@@ 1.1 Background and Motivation
Recently, multimodal recommender systems (MRSs) have garnered significant attention due to their ability to accommodate diverse item information from multiple modalities for enhanced recommendation performance. Compared to canonical recommender systems ranjbar2015imputation; DBLP:journals/eaai/ValcarceLPB19; DBLP:conf/sigir/0001DWLZ020 with no item features (referred to as single-modal recommender systems), MRSs can capture and leverage precise item information (e.g., textual and/or visual features) for recommendations, thereby enhancing the overall capabilities of recommender systems DBLP:conf/aaai/HeM16; DBLP:conf/mm/WeiWN0C20. Notably, while single-modal recommender systems often struggle with sparsity of user–item interactions, MRSs can overcome this limitation by utilizing multimodal features of items. Due to these advantages, MRSs DBLP:conf/mm/Zhang00WWW21; DBLP:conf/www/ZhouZLZMWYJ23; DBLP:conf/mm/ZhouS23; DBLP:conf/mm/Yu0LB23 are shown to substantially outperform single-modal recommendation methods based on collaborative filtering (CF) that rely solely on historical user–item interactions.
@@
Various MRSs have been developed to improve recommendation performance. In particular, thanks to the expressive capability via message passing in graph convolutional networks (GCNs) DBLP:conf/iclr/KipfW17, attention has been paid to GCN-based MRSs DBLP:conf/mm/WeiWN0C20; DBLP:conf/mm/Zhang00WWW21; DBLP:conf/www/ZhouZLZMWYJ23; DBLP:conf/mm/ZhouS23; DBLP:conf/mm/Yu0LB23. For example, prior studies on MRSs learned GCNs separately to process different modalities, distinct from user–item interactions DBLP:conf/mm/WeiWN0C20; DBLP:conf/mm/Zhang00WWW21; DBLP:conf/wsdm/KimKSK24, or constructed an item–item similarity graph based on multimodal information (i.e., multimodal features of items) and then applied GCNs to process the similarity graph DBLP:conf/mm/Zhang00WWW21.
@@
Figure 1: Training time comparison of two GCN-based MRSs under different degrees of modality information on the Baby dataset.
@@
On one hand, user preferences tend to shift quickly under the influence of trends, personal situations, and exposure to new content ju2015using; DBLP:journals/ml/PereiraGAO18; hence, recommender systems need to be flexible to adapt to such dynamic preferences. Especially in environments where the training and inference speed is crucial, model runtime can become a significant bottleneck. In this context, incorporating additional information (i.e., multimodal features of items) into modeling significantly increases computational overhead for GCN-based MRSs. Consequently, MRSs learn multimodal information through GCNs, naturally leading to escalation of training time for GCNs with the increased number of modalities. Figure 1 illustrates how the training time (in seconds) of various MRSs behaves according to different degrees of modality information on the Baby dataset (one of well-known benchmark datasets for MRSs). As shown in Figure 1, incorporating all modalities substantially increases the training time of GCN-based MRSs such as GRCN DBLP:conf/mm/WeiWN0C20 and LATTICE DBLP:conf/mm/Zhang00WWW21, highlighting the heavy computational burden of training-based MRSs. Thus, although GCN-based MRSs (e.g., DBLP:conf/mm/WeiWN0C20, DBLP:conf/mm/Zhang00WWW21) are promising, their computational cost becomes prohibitive in scenarios where models must be updated frequently to track rapidly evolving user preferences.
@@
On the other hand, recent studies on single-modal recommender systems DBLP:conf/cikm/ShenWZSZLL21; DBLP:conf/www/XiaLGLLG22; DBLP:conf/www/LiuLGL0SG23; DBLP:conf/sigir/ParkSS24 have adopted the training-free graph filtering (GF) (also known as graph signal processing) mechanism due to its simplicity and effectiveness. As training-free designs, GF bypasses iterative gradient-based training and instead performs graph filtering along with similarity graphs, thereby resulting in substantially lower computational overhead and much faster inference for recommendation than the case of conventional deep or graph neural network (GNN)-based models. Such GF is a fundamental operation enabling the manipulation of signals defined over the nodes based on the underlying graph topology DBLP:conf/nips/DefferrardBV16; DBLP:journals/tsp/IsufiGSS24. A pioneering study on training-free GF is GF-CF DBLP:conf/cikm/ShenWZSZLL21, which presented a novel approach to constructing an item–item similarity graph and applying GF that does not necessitate model training to enhance recommendation performance. Additionally, recent studies DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25 proposed a normalization scheme with tunable user–item influence and smoothing strength, combined with polynomial graph filters that avoid explicit eigendecomposition.
@@
Despite these advances, the GF-based recommendation methods in DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25 present a key limitation from the perspective of frequency analysis. Normalization techniques in DBLP:conf/cikm/ShenWZSZLL21; DBLP:conf/www/XiaLGLLG22; DBLP:conf/www/LiuLGL0SG23 ensure that all eigenvalues of the underlying item–item similarity graph lie within the interval $[0,1]$, providing a stable theoretical foundation for designing polynomial filters in the frequency domain. In contrast, the normalization and additional adjustment process in DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25 does not enforce this spectral constraint, allowing eigenvalues to extend beyond the unit interval, which distorts the intended frequency response. This effect is magnified on multimodal information due to heterogeneous characteristics, resulting in greater spectral instability than the case of user–item interactions. Additionally, such an effect can be amplified with higher-order polynomials, disproportionately emphasizing components tied to large-magnitude eigenvalues and further degrading frequency control. The aggregation of multiple out-of-band components generates spurious, highly distorted signals. These limitations constitute a fundamental theoretical flaw that calls for a principled reformulation of polynomial graph filtering in recommendation. Moreover, their polynomial filters in DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25 were analytically derived from a fixed frequency response function, which lacks flexibility. This limitation underscores the need for a more adjustable and theoretically grounded filtering strategy for recommender systems.
@@ 1.3 Organization and Notations
$\lambda_{\min}$ Minimum eigenvalue of $\Lambda$
$\lambda_{\max}$ Maximum eigenvalue of $\Lambda$
@@ 2 Related Work
GF-based recommendation. In the context of GF, GCN DBLP:conf/iclr/KipfW17 can be viewed as a parameterized convolutional filter for graphs. A notable GCN-based recommendation method is NGCF DBLP:conf/sigir/Wang0WFC19, which was proposed to learn suitable low-pass filters (LPFs) while capturing high-order collaborative signals present in user–item interactions. LightGCN DBLP:conf/sigir/0001DWLZ020 further demonstrated strong performance by simplifying NGCF, removing both linear transformations and non-linear activations from its GCN layers. However, they are all training-based models that rely on iterative gradient-based optimization over multiple layers, which leads to non-trivial training time and substantial GPU memory consumption, potentially limiting their scalability and ease of deployment in large-scale recommender systems as well as their practicality in latency-sensitive or frequently updated recommendation scenarios.
By bridging the gap between LightGCN and GF approaches, GF-CF DBLP:conf/cikm/ShenWZSZLL21 was introduced, offering satisfactory recommendation accuracy with minimal computational costs due to its training-free design and a closed-form solution for the infinite-dimensional LightGCN. As a follow-up study, PGSP DBLP:conf/www/LiuLGL0SG23 employed a mixed-frequency filter that integrates a linear LPF with an ideal LPF. Nevertheless, both GF-CF and PGSP essentially rely on combinations of linear filters and ideal LPFs in the spectral domain, which restricts the flexibility of their frequency responses. In particular, due to the fact that their filters are restricted to either linear or ideal low-pass forms, once a cutoff frequency is chosen, all spectral components above it are uniformly attenuated, leaving limited flexibility to treat different frequency ranges in a more fine-grained manner. In addition, since the ideal LPF component is defined in terms of the eigenvalues and eigenvectors of the graph operator, its construction relies on (exact or approximate) spectral decomposition, which can become a computational bottleneck when the item–item similarity graph is large.
Turbo-CF DBLP:conf/sigir/ParkSS24 takes a further step toward efficiency by designing polynomial LPFs that retain low-frequency signals without explicitly using an ideal LPF that requires costly matrix decompositions. While this matrix-decomposition-free design substantially reduces runtime, the polynomial filters in Turbo-CF are applied to graph operators whose eigenvalues are not guaranteed to lie within a bounded interval (e.g., $[0,1]$). This mismatch can break the standard assumptions of polynomial graph filtering, making the resulting frequency response harder to analyze and potentially leading to undesirable amplification or attenuation of certain spectral components. In contrast, our proposed MM-GF follows the training-free GF paradigm while explicitly rescaling the graph spectrum so that the eigenvalues fall into a bounded range suitable for polynomial graph filtering. This enables a theoretically grounded design of the filter, preserves desirable low-frequency information, and yields a stable frequency response for multimodal recommendation.
Multimodal recommendation. Studies on MRSs have been actively conducted to boost the performance of recommender systems by leveraging multimodal features of items (i.e., textual and visual features). Early work such as VBPR DBLP:conf/aaai/HeM16 extended the Bayesian personalized ranking (BPR) loss DBLP:journals/corr/abs-1205-2618 by incorporating visual features into matrix factorization, which effectively alleviates item cold-start issues but still models only low-order user–item interactions within a single visual modality. With the rise of graph-based recommendation, GCN-based multimodal models have been proposed to capture high-order connectivity on user–item graphs. For example, GRCN DBLP:conf/mm/WeiWN0C20 refined the user–item bipartite graph by pruning false-positive edges based on multimodal content, thereby improving robustness to noisy implicit feedback at the cost of an additional graph-refining module and iterative training. LATTICE DBLP:conf/mm/Zhang00WWW21 explicitly constructed modality-aware item–item graphs and applied graph convolution to capture high-order semantic item relations, which yields strong accuracy but incurs substantial memory and time overhead because the latent item–item graphs scale quadratically with the number of items.
More recent studies introduced self-supervised and structure-regularized graph learning for multimodal recommendation. BM3 DBLP:conf/www/ZhouZLZMWYJ23 bootstrapped latent representations from user–item interactions and multimodal features using a contrastive self-supervised objective, avoiding negative sampling and auxiliary graphs while still requiring the optimization of a deep GCN backbone and multiple multimodal losses. FREEDOM DBLP:conf/mm/ZhouS23 simplified LATTICE by freezing the latent item–item graph and incorporating a degree-sensitive edge-pruning strategy on the user–item graph, which improves both recommendation accuracy and memory efficiency on large graphs; however, it remains a fully trainable GNN-based architecture with non-trivial hyperparameter sensitivity (e.g., pruning ratio, loss weights). In MGCN DBLP:conf/mm/Yu0LB23, noisy modality features were explicitly purified and enriched in separate graph views before being fused by a behavior-aware module, improving robustness to noisy content at the cost of additional denoising networks and fusion components. Beyond standard bipartite graphs, LGMRec DBLP:conf/aaai/GuoL0WSR24 constructed a modality-aware hypergraph to jointly model local and global user interests, while PGL DBLP:conf/aaai/Yu0LB25 extracted principal subgraphs from user–item interactions to preserve individual (high-frequency) structural information. These models enhance expressiveness by introducing hypergraph modules and principal subgraph learning, but their complex architectures and training objectives further increase computational and implementation overhead.
In summary, existing multimodal recommenders can be grouped into 1) shallow CF models with side multimodal features, and 2) training-based graph models that learn high-capacity GNNs, auxiliary graphs, or self-supervised objectives. While the latter achieve strong accuracy, they typically require expensive end-to-end training, careful negative sampling or contrastive loss design, and additional graph-learning modules. In contrast, our proposed MM-GF takes a complementary, training-free approach: it models multimodal recommendation as spectral polynomial graph filtering, using closed-form filters defined on preconstructed user–item and item–item graphs. By avoiding parameterized GNN layers and negative-sampling-based optimization, MM-GF maintains the ability to exploit high-order collaborative and multimodal signals, while enjoying lower computational complexity and a simpler, model-agnostic deployment pipeline.
@@ 3.1 Notion of GF
We provide fundamental principles of GF (or equivalently, graph signal processing) DBLP:journals/pieee/OrtegaFKMV18; DBLP:conf/cikm/ShenWZSZLL21; DBLP:conf/sigir/ParkSS24. First, we consider an undirected graph $\mathcal{G=(V,E)}$, which is represented by an adjacency matrix $A$ that indicates the presence or absence of edges between nodes. The Laplacian matrix $L$ of $\mathcal{G}$ is defined as L=DAL=D-A, where $D$ is the degree matrix of $A$ DBLP:journals/siamdm/GroneM94. Meanwhile, a graph signal on $\mathcal{G}$ is expressed as $\mathbf{x}\in\mathbb{R}^{|\mathcal{V}|}$, where each $x_{i}$ indicates the signal strength at the corresponding node $i$ in $\bf x$}. The smoothness of a graph signal $\mathbf{x}$ can be mathematically measured by
@@ 4 Methodology
In this section, we elaborate on the proposed MM-GF method and its computational complexity as well as our research motivation.
@@ 4.1 Motivation and Challenges
Conventional training-free GF-based recommendation methods DBLP:conf/cikm/ShenWZSZLL21; DBLP:conf/www/LiuLGL0SG23 start by constructing a graph structure, where nodes represent items and edges correspond to item–item similarities. The graph construction process is formulated as follows:
@@
Recently, DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25 utilized asymmetric normalization on $R$ to regularize the popularity of users or items before calculating $\tilde{P}$, which is formulated as follows:
@@
However, as critical issues, introducing two hyperparameters (i.e., $\alpha$ and $s$) may alter the spectral characteristics of the underlying graph. Specifically, the work in DBLP:conf/cikm/ShenWZSZLL21 proved that all the eigenvalues of $\tilde{P}=\tilde{R}^{T}\tilde{R}$ where $\tilde{R}=D_{r}^{-0.5}RD_{c}^{-0.5}$ are in $[0,1]$. Nonetheless, as long as parameters $\alpha$ and $s$ are concerned, one cannot always assure that $\bar{P}$ in Eqs. (5) and (6) satisfies the aforementioned condition in Section 1.1 (i.e., the unit interval of eigenvalues) when $\alpha\neq 0.5$ and/or $s\neq 1$. In Figure 2, we present two cases: Case 1) $\alpha=0.5$ and $s=1$, corresponding to symmetric normalization, and Case 2) $\alpha=0.7$ and $s=0.6$, corresponding to asymmetric normalization with an additional adjustment process. The observations from this figure pose practical challenges below.
@@
As observed in Figure 2, when we construct the graph operators exactly as in previous GF-based methods DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25, the resulting eigenvalue spectrum deviates substantially from the canonical interval $[0,1]$, with plenty of eigenvalues lying far below 0 or far above 1; this mismatch becomes problematic precisely when the polynomial LPFs from prior studies are naïvely applied to such spectra, since they were originally derived under the assumption of eigenvalues confined to $[0,1]$. Symmetric normalization in DBLP:conf/cikm/ShenWZSZLL21; DBLP:conf/www/LiuLGL0SG23; DBLP:conf/www/XiaLGLLG22 typically follow that the spectrum lies within $[0,1]$, which was proved by DBLP:conf/cikm/ShenWZSZLL21. When eigenvalues are mapped out of this range, the polynomial filter may yield inaccurate frequency responses, deviating from its designed frequency characteristics. Moreover, as the polynomial order increases, the filter response for large-magnitude eigenvalues can grow dramatically; this results in a disproportionate emphasis on certain frequency components, thereby degrading the model performance. For example, consider a polynomial LPF $h(\lambda)=1-\lambda^{2}$. In a bounded frequency range of $[0,1]$, filtered values close to 0 correspond to high-frequency components, whereas those close to 1 correspond to low-frequency components. However, if the spectrum contains eigenvalues such as $\lambda_{1}\simeq-1$ and $\lambda_{2}\simeq 14$, as shown in Figure 2(b), then their corresponding filter responses become $h(\lambda_{1})\simeq 0$ and $h(\lambda_{2})\simeq-195$, respectively. This phenomenon results in signals being assigned to anomalous frequencies instead of the targeted ones. Furthermore, this problem becomes particularly critical for multimodal information, where signals are more susceptible to distortion compared to the case of user–item interactions. For example, if the spectrum contains an eigenvalue $\lambda_{3}\simeq 70$, as illustrated in Figures 2(d) and 2(f), then its corresponding filter response becomes $h(\lambda_{3})\simeq-4899$, leading to far more severe distortion than the case of user–item interactions. Ultimately, since MRSs require integrating signals across multiple modalities, aggregating such severely distorted signals can undermine the overall reliability of the recommendation process.
Additionally, despite the extensive literature on CF, there currently exists no training-free popularity-balancing formulation that preserves the spectral properties required for GF. Prior approaches addressing popularity bias (e.g., DBLP:conf/sigir/0002YKL25; DBLP:journals/tors/PengSLM25) rely on learning-based adjustments and therefore cannot be applied when an analytically grounded, training-free mechanism is desired. This limitation further motivates the need for a principled normalization strategy, which our framework is designed to provide.
@@
Addressing a different issue, while recent GF models DBLP:conf/sigir/ParkSS24; DBLP:conf/www/ParkYS25; DBLP:conf/www/KimCPS25 demonstrated that a polynomial graph filter can be implemented as the form of $\sum_{k=1}^{K}a_{k}\bar{P}^{k}$, their empirical application was confined only to specific filters. This limited scope highlights a gap in the broader applicability of the GF methodology. Moreover, it is of paramount importance to effectively aggregate information across multiple modalities for accurate multimodal recommendations. In light of these challenges, a key question arises: "How can we design an efficient and effective GF method for multimodal recommendations by not only addressing the issue of filter design but also maximally exploiting heterogeneous characteristics across modalities?" To answer this question, we will outline the proposed MM-GF method tailored for multimodal recommendations in the following subsection.
Figure 3: The schematic overview of MM-GF.